# Lower Bound of Nodal Sets in Elliptic Homogenization and Functions with Strong Maximum Principle

### v1

@@ Abstract
In this note, we first try to prove a uniform lower bound of nodal volume in elliptic homogenization setting. This lower bound is far from optimal. But, we can prove a constant lower bound in dimension two. Motivated by the proof, we extend this results to more general settings. To be more specific, we prove that the nodal volume has a constant lower bound for all continuous functions with strong maximum princple. Our result works for general functions beyond solutions to elliptic PDEs.
@@ 1 Introduction
The nodal sets of elliptic PDEs is a challenging problem in analysis. In the work of Lin [MR1090434], it has been proved that the nodal sets of elliptic PDEs has Hausdorff measure less than or equal to n1n-1. Then, it is natural to ask whether the nodal volume has upper bound and lower bound locally and globally. When considering the Laplace eigenfunction on Riemannian manifold, Yau made the following conjecture:
@@
Let $(M^{n},g)$ be a closed manifold and $u$ be a nonconstant eigenfunction that satifies $\Delta_{g}u+\lambda u=0$ with $\lambda>0$. Then
@@
For the lower bound, it can be reduced to Nadirashvili’s conjecture. It was conjectured in [nadirashvilli1997geometry] that, if a harmonic function vanishes at zero point, then the Hausdorff measure of nodal sets has a constant lower bound in unit ball. This conjecture has been solved by Logunov in [MR3739232]. Another version of Nadirashvili’s conjecture was solved by the authors and Wang in [li2025nadirashviliconjectureellipticpdes]. Actually, in [MR4814921], Logunov, Lakshmi Priya and Sartori proved a much stronger version. They proved an almost sharp lower bound of nodal volume of elliptic PDEs. We will recall the main theorem here:
@@
Indeed, this lower bound works for more genenral elliptic PDEs with smooth and $C^{1}$ coefficients. In their work, they made the following conjecture:
@@
However, this conjecture still remain open. Besides the estimate of lower bound, there are also many results concerning the upper bound of nodal sets for elliptic PDEs and especially the eigenfunction. In [MR3739231], Logunov proved a polynomial upper bound of nodal sets of Laplace eigenfunction on smooth manifold. Before that, Donnelly and Fefferman proved the Yau’s conjecture completely for Laplace eigenfunction on real-analytic manifold in [MR943927] and [MR1035413]. For more general settings, Lin proved the sharp upper bound of nodal volume for elliptic PDEs with real-analytic coefficients in [MR1090434]. Hardt and Simon proved a exponential upper bound for elliptic PDEs with Lipschitz coefficient in [MR1010169]. Later, Han and Lin gave a different proof in [Hanlin94]. Han also used Schauder estimate to stratify the nodal sets in [MR1794573]. In [MR3688031], Naber and Valtorta gave a upper bound estimate for elliptic PDEs with Lipschitz coefficients. There are also many results concerning solutions to other equations including parabolic equations, bi-harmonic functions and so on. The reader may refer [MR3928756][huang2024nodal][huang2023volume] and other related references.
An important theory in elliptic PDE is the homogenization theory. There is a small parameter $\varepsilon$ in the setting. The coefficients are highly oscalliting which makes previous results failing. In [MR3952693], Lin and Shen first proved a uniform upper bound of nodal volume in elliptic homogenization. They used the compactness method to proved the doubling inequality. Later, Kenig, Zhu and Zhuge gave a explicit formula of upper bound in [MR4387203]. Also, this upper bound is far from optimal. But they used conformal mapping to proved a refined upper bound in dimension two. In [MR4764741], Lin and Shen also studied the upper bound of critical sets in elliptic homogenization. They used the method of Naber and Valtorta in their work [MR3688031]. They also proved the upper bound of critical sets in dimension two with a different method in [MR4669024].
It is a natural question to ask whether we can prove a similar lower bound of nodal volume in homogenization. Let us start with basic settings. In the first part of our note, we will consider the following elliptic PDE in $B(0,1)\subset\mathbb{R}^{n}$:
@@
We are unlikely to prove the first version of results. So in this note, we try to prove a lower bound depending on the doubling index when $\varepsilon>0$ is small. Also, we hope we can get a explicit formula of the lower bound for all $\varepsilon\in(0,1)$. Before that, we will first introduce the definition of doubling index.
@@
with $B(x,r)\subset B(0,1)$. Sometimes, we will write $N(B)=N(x,r)$ if $B=B(x,r)$. Also we will simply write $N(r)$ if $x$ is fixed or $x$ is the zero point. If we want to specify the choice of function, we will write $N_{u}(x,r)$.
@@
Assume $A\in\mathcal{A}$. Let $u_{\varepsilon}\in H^{1}(B(0,1))$ be a solution of $\mathcal{L}_{\varepsilon}(u_{\varepsilon})=0$ in $B(0,1)\subset\mathbb{R}^{n},n\geq 3$, where $0<\varepsilon<1$. Let $N_{0}=N_{u_{\varepsilon}}(0,1)$. Then there exsits a constant $\varepsilon_{0}$ depending on $A,n,N_{0}$ such that the following estimate holds: if $0<\varepsilon<\varepsilon_{0}$
@@
whenever $u_{\varepsilon}(0)=0$ and $C$ only depends on $A,n$.
Combining Theorem 1.2, we can conclude the following uniform lower bound estimate of homogenization equtions.
@@
In dimension two, we use a different approach. We will only use blow up argument and maximum principle to derive a constant lower bound.
@@
Assume $A\in\mathcal{A}$. Let $u_{\varepsilon}\in H^{1}(B(0,1))$ be a solution of $\mathcal{L}_{\varepsilon}(u_{\varepsilon})=0$ in $B(0,1)\subset\mathbb{R}^{2}$. Then there exsits a constant $C_{0}>0$ depending on $A$ only such that the following estimate holds
@@
Motivated by the proof of above theorem, we try to give a much simpler proof of Nadirashvili’s conjecture in dimension two. Indeed, we can prove Nadirashvili’s conjecture in a more general settings. We only assume the function satisfies strong maximum principle and is continuous. Here, we recall the content of Nadirashvili’s conjecture for harmonic functions.
@@
We mention again that this theorem actually works for a wide range of solutions of PDEs not only solutions to elliptic PDEs. And the lower bound $2$ is sharp in dimension two. In fact, if we assume the set $\left\{u=0\right\}\cap B_{1}$ is countably rectifiable, then this theorem is easy to prove by coarea formula. Here we do not assume it is countably rectifiable. So we need a new analytical method to analyze the lower bound of the nodal set.
Notations. We use the following notations in this article.
  • Throughout this article, we denote positive constants by $C,C_{1},\cdots$. We may write $C(a_{1},a_{2},\cdots)$ to highlight its dependence on parameters $a_{1},a_{2},\cdots$. The value of $C$ may vary from line to line.
  • In this article, we denote the ball centered at $x$ with radius $r$ by $B(x,r)=\{y\in\mathbb{R}^{n}:|x-y|\leq r\}$. We will write $B_{r}$ if $x$ is fixed or it is 0.
@@
Acknowledgements. Jiahuan Li was supported by National Natural Science Foundation of China [grant number 12141105]. The authors would like to thank anyone for reading the draft of this note and useful comments.
2 Unifrom Lower Bound for Elliptic Homogenization
2.1 Property from Homogenization Theory
In this subsection, we will first collect several important properties from homogenization theory. Most of the theorem can be found in [MR4764741] and [MR4387203].
@@ 2.2 Approximation and Volume Estimate
As a result, we can get a lower bound of $k^{+}$ as the following
@@
So we know that there are at least $\frac{k}{3}$ index that are both $S^{+}$-good and $S^{-}$-good. For any $i$ that is both $S^{+}$-good and $S^{-}$-good, we assume $u_{0}(x_{1})=\sup_{B_{i}}u_{0}$. By standard gradient estimate, we know
@@
With this lemma, we can prove our final results.
@@
Indeed, we just need to verify conditions in Lemma 2.4. The second condition is satisfied by Theorem 2.2 after scaling and doubling inequality. We just need to verify the first condition and the doubling inequality. Let us start with the following estimate
@@ 3 Special Case in Two Dimension
In this section, we will try to prove the constant lower bound in dimension two. We will prove a stronger conclusion. And we will use methods from elliptic PDEs. All the functions in this sections are defined in $\mathbb{R}^{2}$. First, we need the following constant lower bound for elliptic PDEs. This lemma does not imply Theorem 1.7 due to the existence of $\varepsilon$.
@@
$\mathrm{div}\left(A\nabla u\right)=0\ in\ Q_{k}.$ (3.2)
We can assume that $u$ is not always zero, otherwise the original theorem would obviously hold. We assume that
$M_{s}:=\underset{Q_{s}}{\sup}u,\ m_{s}:=\underset{Q_{s}}{\inf}u\ \forall s>0.$ (3.3)
@@
Hence by the continuity of function $u$, we know that for any $s>0$, there exists $x_{s}\in\partial Q_{s}$ such that $u\left(x_{s}\right)=0$. Therefore, for any positive integer $t\leq k$, we can find a closed cube $Q^{t}$ whose side longth is 2, such that the center of $Q^{t}$ is an integer point and $\left\{u=0\right\}\cap\frac{1}{2}Q^{t}\neq\emptyset$. By Lemma 3.1 and the periodicity of $A$ we know there exists a constant $C$ depending only on $A$ such that
$\mathcal{H}^{1}(\left\{u=0\right\}\cap Q^{t})\geqslant C.$ (3.5)
The constant is independent of $x$ by periodicity. We define the number of intersections between cubes as follows
$b_{j}:=\#\left\{i|Q^{i}\cap Q^{j}\neq\emptyset\right\}.$ (3.6)
@@
So we can find $k/C_{1}$ non-intersecting cubes from $\left\{Q^{t}\right\}_{t=1}^{k}$. So there exists a constant $C$ only depending on $A$ such that
@@
In fact, We can prove the following more general conclusion.
@@
We assume that $u$ is a solution to the equation $\mathrm{div}\left(A\nabla u\right)=0$ in $B_{1}$ and the matrix satisfies the following conditions.
@@
for some $\lambda>0$ .Then there exists a constant $C$ depending only on $\lambda$ such that if $u(0)=0$ then
$\mathcal{H}^{1}(\left\{u=0\right\}\cap B_{1})\geq C.$ (3.10)
This theorem implies that the lower bound is independent of Lipschitz constant of coefficients. To prove this theorem, we need two classic results and two important lemmas.
@@
Assume $A$ and $u$ satisfies the assumption in Theorem 2.3, then there exists a constant $C$ denpending only on $\lambda$ such that if $u(0)\geq 0$ then
@@
For convenience, we will first use the following notation.
@@
We need the following classical Schauder estimate.
@@
Then we will prove the following doubling inequality.
@@
We use the method of contradiction to prove this lemma. If this lemma is not holds, then we can find $\left\{A_{k}\right\}_{k=1}^{\infty}\subset A\left(\lambda,M\right),\left\{u_{k}\right\}_{k=1}^{\infty}$ such that $\mathrm{div}\left(A_{k}\nabla u_{k}\right)=0\ in\ B_{1},u_{k}\left(0\right)=0$, $u_{k}$ is not always zero and
@@
By scaling we can assume $\underset{B_{2/5}}{\sup}|u_{k}|=1$, hence $\underset{B_{4/5}}{\sup}|u_{k}|\leq 1+\frac{1}{k}$. By Schauder estimate and taking subsequence, we can assume there exists $u^{*}\in C^{1,\alpha}\left(B_{3/4}\right)$ such that
@@
And by taking subsequence we also can assume that there exists $A^{*}\in A\left(\lambda,M\right)$ such that
@@
Hence we know $u^{*}$ is the weak solution of $\mathrm{div}\left(A^{*}\nabla u\right)=0\ in\,\,B_{3/4}$, and by assumption of contradiction we know
@@
With above lemmas, we can prove the weakened version of Theorem 3.2. We will later use this results to prove the stronger version.
@@
Assume $A\in{A}\left(\lambda,M\right)$, $\mathrm{div}\left(A\nabla u\right)=0\ in\ B_{1}$ ,$u(0)=0$ and $u$ is not zero. There exists a constant $C$ depending only on $\lambda,M$ such that
@@
Divide interval $\frac{1}{2},\frac{7}{10}]$ equally into
@@
WLOG, we assume that $\underset{B_{2/5}}{\sup}|u|=1$. Then by Lemma 3.4 , we can asuume
@@
Hence by Lemma 3.7, there exists a constant $s_{0}=s_{0}(\lambda,M)$ such that if $s\leq s_{0}$ than
@@
Fix $s=s_{0}$ and take $k=3\left(\log_{s_{0}}2\right)C_{3}\left(\lambda,M\right)N$. Then, we have the estimate of $k^{+}$
@@
By scaling, Schauder estimates and Lemma 3.7 we know
@@
Hence, we can estimate $u$ as below
@@
Finally, we can prove the general conclusion.
@@
Also, by similar argument of Lemma 2.4, we can prove the following theorem.
@@
Assume that $A\in{A}\left(\lambda,M\right)$, $\mathrm{div}\left(A\nabla u_{0}\right)=0\ in\ B_{1}$ $u_{0}(0)=0$ and $u_{0}$ is not zero. The function $u$ is continuous. There exists two constant $\varepsilon,C$ depending only on $\lambda,M$ such that if
@@
It tells us that the nodal volume still has lower bound under small perturbation. But the quantitative lower bound is still unknow.

### v2

@@ Abstract
In this note, we first prove a uniform lower bound of nodal volume in elliptic homogenization setting. This lower bound is far from optimal. But, we can prove a constant lower bound in dimension two. Motivated by the proof, we extend this result to more general settings. To be more specific, we prove that the nodal volume has a constant lower bound for all continuous functions with strong maximum principle. Our results work for general functions beyond solutions to elliptic PDEs.
@@ 1 Introduction
The analysis of nodal sets in elliptic PDEs is a challenging problem. In the work of Lin [MR1090434], it has been proved that the nodal sets of elliptic PDEs have Hausdorff measure less than or equal to n1n-1. Then, it is natural to ask whether the nodal volume has upper bound and lower bound locally and globally. When considering the Laplace eigenfunction on Riemannian manifold, Yau made the following conjecture:
@@
Let $(M^{n},g)$ be a closed and compact manifold and $u$ be a nonconstant eigenfunction that satisfies $\Delta_{g}u+\lambda u=0$ with $\lambda>0$. Then
@@
For the lower bound, it can be reduced to Nadirashvili’s conjecture. It was conjectured in [nadirashvilli1997geometry] that, if a harmonic function vanishes at origin, then the Hausdorff measure of nodal sets has a constant lower bound in unit ball. This conjecture has been solved by Logunov in [MR3739232]. Another version of Nadirashvili’s conjecture was solved by the authors and Wang in [li2025nadirashviliconjectureellipticpdes]. Actually, in [MR4814921], Logunov, Lakshmi Priya and Sartori proved a much stronger version. They proved an almost sharp lower bound of nodal volume of elliptic PDEs. We will recall the main theorem here:
@@
Indeed, this lower bound works for more genenral elliptic PDEs with $C^{1}$ coefficients. In their work, they made the following conjecture:
@@
However, this conjecture remains open. Besides the estimate of lower bound, there are also many results regarding the upper bound of nodal sets for elliptic PDEs and especially the Laplace eigenfunction. In [MR3739231], Logunov proved a polynomial upper bound of nodal sets of Laplace eigenfunction on smooth manifold. Before that, Donnelly and Fefferman proved the Yau’s conjecture completely for Laplace eigenfunction on real-analytic manifold in [MR943927] and [MR1035413]. For more general settings, Lin proved the sharp upper bound of nodal volume for elliptic PDEs with real-analytic coefficients in [MR1090434]. Hardt and Simon proved a exponential upper bound for elliptic PDEs with Lipschitz coefficient in [MR1010169]. Later, Han and Lin gave a different proof in [Hanlin94]. Han also used Schauder estimate to stratify the nodal sets in [MR1794573]. In [MR3688031], Naber and Valtorta gave a upper bound estimate for elliptic PDEs with Lipschitz coefficients. There are also many results concerning solutions to other equations including parabolic equations, bi-harmonic functions and so on. The reader may refer [MR3928756][huang2024nodal][huang2023volume] and other related references.
An important theory in elliptic PDE is the homogenization theory. There is a small parameter $\varepsilon$ in the setting. The coefficients are highly oscalliting which makes previous results fail to apply. In [MR3952693], Lin and Shen first proved a uniform upper bound of nodal volume in elliptic homogenization. They used the compactness method to proved the doubling inequality. Later, Kenig, Zhu and Zhuge gave an explicit formula of upper bound in [MR4387203]. Also, this upper bound is far from optimal. But they used conformal mapping to prove a refined upper bound in dimension two. In [MR4764741], Lin and Shen also studied the upper bound of critical sets in elliptic homogenization. They used the method of Naber and Valtorta in their work [MR3688031]. They also proved the upper bound of critical sets in dimension two with a different method in [MR4669024].
It is a natural question to ask whether we can prove a similar lower bound of nodal volume in homogenization. We will begin with basic settings. In the first part of our note, we will consider the following elliptic PDE in $B(0,1)\subset\mathbb{R}^{n}$:
@@
We do not expect to prove the first version of results. So in this note, we try to prove a lower bound depending on the doubling index when $\varepsilon>0$ is small. Also, we hope we can get a explicit form of the lower bound for all $\varepsilon\in(0,1)$. Before that, we will first introduce the definition of doubling index.
@@
with $B(x,r)\subset B(0,1)$. Sometimes, we will write $N(B)=N(x,r)$ if $B=B(x,r)$. Also we will simply write $N(r)$ if $x$ is fixed or $x$ is the origin. If we want to specify the choice of function, we will write $N_{u}(x,r)$.
@@
Assume $A\in\mathcal{A}$. Let $u_{\varepsilon}\in H^{1}(B(0,1))$ be a solution of $\mathcal{L}_{\varepsilon}(u_{\varepsilon})=0$ in $B(0,1)\subset\mathbb{R}^{n},n\geq 3$, where $0<\varepsilon<1$. Let $N_{0}=N_{u_{\varepsilon}}(0,1)$. Then there exists a constant $\varepsilon_{0}$ depending on $A,n,N_{0}$ such that if $0<\varepsilon<\varepsilon_{0}$, the following estimate holds:
@@
whenever $u_{\varepsilon}(0)=0$ and $C$ only depends on $A$ and $n$.
Combining Theorem 1.2, we conclude the following uniform lower bound estimate of homogenization equations:
@@
In dimension two, we use a different approach. We will only use blow up argument and maximum principle to derive a constant lower bound:
@@
Assume $A\in\mathcal{A}$. Let $u_{\varepsilon}\in H^{1}(B(0,1))$ be a solution of $\mathcal{L}_{\varepsilon}(u_{\varepsilon})=0$ in $B(0,1)\subset\mathbb{R}^{2}$. Then there exsits a constant $C_{0}>0$ depending only on $A$ such that the following estimate holds
@@
Motivated by the proof of above theorem, we try to give a much simpler proof of Nadirashvili’s conjecture in dimension two. Indeed, we can prove Nadirashvili’s conjecture in a more general setting. We only assume the function satisfies strong maximum principle and is continuous. Here, we recall the content of Nadirashvili’s conjecture for harmonic functions:
@@
We mention again that this theorem actually works for a wide range of solutions of PDEs not only solutions to elliptic PDEs. This results shows that the lower bound in dimension two is a conseqence of maximum principle rather than the structure of PDE. The lower bound $2$ is sharp. In fact, if we assume the set $\left\{u=0\right\}\cap B_{1}$ is countably rectifiable, then this theorem is easy to prove by coarea formula. Here we do not assume it is countably rectifiable. Thus, we need a new analytical method to prove the lower bound of the nodal set.
Notations. We use the following notations in this note.
  • We denote positive constants by $C,C_{1},\cdots$. We may write $C(a_{1},a_{2},\cdots)$ to highlight its dependence on parameters $a_{1},a_{2},\cdots$. The value of $C$ may vary from line to line.
  • We denote the ball centered at $x$ with radius $r$ by $B(x,r)=\{y\in\mathbb{R}^{n}:|x-y|\leq r\}$. We will write $B_{r}$ if $x$ is fixed or it is 0.
@@
Acknowledgements. Jiahuan Li was supported by National Natural Science Foundation of China [grant number 12141105]. The authors would like to thank anyone who read the draft of this note and give useful comments.
2 Uniform Lower Bound for Elliptic Homogenization
2.1 Properties from Homogenization Theory
In this subsection, we will first collect several important properties from homogenization theory. Most of these results can be found in [MR4764741] and [MR4387203].
@@ 2.2 Approximation and Volume Estimate
As a result, we obtain a lower bound of $k^{+}$ as the following
@@
So we know that there are at least $\frac{k}{3}$ indices that are both $S^{+}$-good and $S^{-}$-good. For any $i$ that is both $S^{+}$-good and $S^{-}$-good, we assume $u_{0}(x_{1})=\sup_{B_{i}}u_{0}$. By standard gradient estimate, we know that
@@
With this lemma, we can prove our final results:
@@
Indeed, we only need to verify conditions in Lemma 2.4. The second condition is satisfied by Theorem 2.2 after scaling and doubling inequality. We just need to verify the first condition and the doubling inequality. Let us start with the following estimate
@@ 3 Special Case in Two Dimension
In this section, we will try to prove the constant lower bound in dimension two. We will prove a stronger conclusion. And we will use methods from elliptic PDEs. All the functions in this sections are defined in $\mathbb{R}^{2}$. First, we need the following constant lower bound for elliptic PDEs. This lemma does not imply Theorem 1.7 due to the presence of $\varepsilon$.
@@
$\mathrm{div}\left(A\nabla u\right)=0\ in\ Q_{k}$ (3.2)
We can assume that $u$ is not identically zero, otherwise the original theorem would obviously hold. We assume that
$M_{s}:=\underset{Q_{s}}{\sup}u,\ m_{s}:=\underset{Q_{s}}{\inf}u\ \forall s>0$ (3.3)
@@
Hence by the continuity of $u$, we know that for any $s>0$, there exists $x_{s}\in\partial Q_{s}$ such that $u\left(x_{s}\right)=0$. Therefore, for any positive integer $t\leq k$, we can find a closed cube $Q^{t}$ whose side length is 2, such that the center of $Q^{t}$ is an integer point and $\left\{u=0\right\}\cap\frac{1}{2}Q^{t}\neq\emptyset$. By Lemma 3.1 and the periodicity of $A$, we know there exists a constant $C$ depending only on $A$ such that
$\mathcal{H}^{1}(\left\{u=0\right\}\cap Q^{t})\geqslant C$ (3.5)
This constant is independent of $x$ by periodicity. We define the number of intersections between cubes as follows
$b_{j}:=\#\left\{i|Q^{i}\cap Q^{j}\neq\emptyset\right\}$ (3.6)
@@
So we can find at least $k/C_{1}$ non-intersecting cubes from $\left\{Q^{t}\right\}_{t=1}^{k}$. So there exists a constant $C$ only depending on $A$ such that
@@
In fact, We can prove a more general conclusion:
@@
We assume that $u$ is a solution to the equation $\mathrm{div}\left(A\nabla u\right)=0$ in $B_{1}$ and the matrix satisfies the following conditions:
@@
for some $\lambda>0$ .Then there exists a constant $C$ depending only on $\lambda$ such that if $u(0)=0$, then
$\mathcal{H}^{1}(\left\{u=0\right\}\cap B_{1})\geq C$ (3.10)
This theorem implies that the lower bound is independent of Lipschitz constant of coefficients. To prove this theorem, we need two classic results and two important lemmas:
@@
Assume $A$ and $u$ satisfy the assumption in Theorem 2.3, then there exists a constant $C$ denpending only on $\lambda$ such that if $u(0)\geq 0$, then
@@
For convenience, in the following argument, we will use the following notation:
@@
We need the following classical Schauder estimate:
@@
Then we will prove the following doubling inequality:
@@
We argue by contradiction. If this lemma does not holds, then we can find $\left\{A_{k}\right\}_{k=1}^{\infty}\subset A\left(\lambda,M\right),\left\{u_{k}\right\}_{k=1}^{\infty}$ such that $\mathrm{div}\left(A_{k}\nabla u_{k}\right)=0\ in\ B_{1},u_{k}\left(0\right)=0$, $u_{k}$ is not identically zero and
@@
By scaling, we can assume $\underset{B_{2/5}}{\sup}|u_{k}|=1$, hence $\underset{B_{4/5}}{\sup}|u_{k}|\leq 1+\frac{1}{k}$. By Schauder estimate and taking subsequence, we can assume that there exists $u^{*}\in C^{1,\alpha}\left(B_{3/4}\right)$ such that
@@
And by taking subsequence, we also can assume that there exists $A^{*}\in A\left(\lambda,M\right)$ such that
@@
Hence we know $u^{*}$ is the weak solution of the equation $\mathrm{div}\left(A^{*}\nabla u\right)=0\ in\,\,B_{3/4}$, and by our assumption, we know
@@
With above lemmas, we can prove the weakened version of Theorem 3.2. We will later use this results to prove the stronger version:
@@
Assume $A\in{A}\left(\lambda,M\right)$, $\mathrm{div}\left(A\nabla u\right)=0\ in\ B_{1}$ ,$u(0)=0$ and $u$ is not zero. Then there exists a constant $C$ depending only on $\lambda,M$ such that
@@
Divide the interval $\frac{1}{2},\frac{7}{10}]$ equally into
@@
WLOG, we assume that $\underset{B_{2/5}}{\sup}|u|=1$. Then by Lemma 3.4 , we can assume
@@
Hence by Lemma 3.7, there exists a constant $s_{0}=s_{0}(\lambda,M)$ such that if $s\leq s_{0}$, then
@@
Take $s=s_{0}$ and $k=3\left(\log_{s_{0}}2\right)C_{3}\left(\lambda,M\right)N$. Then, we have the estimate of $k^{+}$
@@
By scaling, Schauder estimate and Lemma 3.7, we know
@@
Thus, we can estimate $u$ as below
@@
Finally, we can prove the general conclusion:
@@
Also, by similar argument of Lemma 2.4, we can prove the following theorem:
@@
Assume that $A\in{A}\left(\lambda,M\right)$, $\mathrm{div}\left(A\nabla u_{0}\right)=0\ in\ B_{1}$ $u_{0}(0)=0$ and $u_{0}$ is not zero. The function $u$ is continuous. There exists two constant $\varepsilon,C>0$ depending only on $\lambda,M$ such that if
@@
It shows that the nodal volume still has lower bound under small perturbation. But the quantitative lower bound is still unknow.